### v1

# Reward Poisoning Attack Against Offline Reinforcement Learning

Abstract
We study the problem of reward poisoning attacks against general offline reinforcement learning with deep neural networks for function approximation. We consider a black-box threat model where the attacker is completely oblivious to the learning algorithm and its budget is limited by constraining both the amount of corruption at each data point, and the total perturbation. We propose an attack strategy called ‘policy contrast attack’. The high-level idea is to make some low-performing policies appear as high-performing while making high-performing policies appear as low-performing. To the best of our knowledge, we propose the first black-box reward poisoning attack in the general offline RL setting. We provide theoretical insights on the attack design and empirically show that our attack is efficient against current state-of-the-art offline RL algorithms in different kinds of learning datasets.
Machine Learning, ICML
1 Introduction
Reinforcement learning is a popular learning paradigm for studying sequential decision-making problems (Sutton & Barto, 2018; Henderson et al., 2018), which are the essence of a majority of real-world applications. Specifically, reinforcement learning in the offline learning setting is more practical in many scenarios (Levine et al., 2020; Agarwal et al., 2020). Here, the training data is collected in an offline manner. For example, in the medical treatment problem, the offline data can be a recording of how a human doctor treats a patient. In contrast, collecting online data for the medical treatment problem would require the learner to treat patients during training, which is both dangerous and infeasible.
Despite theoretical understanding and empirical success in simulated environments, offline reinforcement learning faces a threat in the real world. In many applications, the offline data is usually based on human feedback (Zheng et al., 2018; Kiran et al., 2021). An adversary can poison the reward signal in the training dataset. When the data is provided by a third party, multiple adversaries can collude to corrupt different parts of the data to achieve the same goal of misleading the learning process. Current state-of-the-art offline reinforcement learning algorithms (Kumar et al., 2020; Fujimoto & Gu, 2021; Kidambi et al., 2020; Wu et al., 2019; Cheng et al., 2022) only consider ideal environments without such threats, which motivates us to investigate the vulnerabilities of these algorithms to such attacks. We focus on developing practical reward-poisoning attacks where the learning algorithm is no longer able to learn a high-performing policy under the attack. In general, a practical attacker should be black-box, that is, it should be oblivious to the learning algorithm and the training process of the learner. The corruption it applies to the reward signal in the dataset should also be limited. We formally introduce our criteria for an attack to be practical in Section 2.
Challenges: We find the following challenges for developing a practical reward poisoning attack:
  1. 1.
    Fixed training dataset: In the online learning setting, an attacker can mislead the learner to collect data that only covers states and actions with low long-term rewards so that the learner never observes how a high-performing policy would work making it fail to learn a high-performing policy (Xu et al., 2022). In contrast, the training dataset is fixed in the offline learning setting. The learner can always observe how an expert would behave in the dataset regardless of the attack. The attacker needs to convince the agent through reward corruption that certain expert behavior is bad even if it frequently occurs in the dataset with high true rewards that indicate its high performance.
  2. 2.
    No access to training process: A majority of online learning attacks need to observe the behavior of the learning algorithm during the training time (Zhang et al., 2020; Rakhsha et al., 2020), which is practical in that case because the learner needs to interact with the online environment. From such observations, the attacker can infer the current status of the learning algorithm, for example, the Q function it learns. However, in the offline setting, there is no interaction between the learner and the environment. The attacker has to infer the status of the learning agent through the training data only.
  3. 3.
    Limited learnable policies: To make offline learning efficient, a learning algorithm needs to be pessimistic (Levine et al., 2020). That is, if the behavior of a policy is barely covered by the dataset, then the algorithms should not output such a policy as it is unable to properly evaluate the policy. As a result, the policies that are not covered by the dataset can be fundamentally impossible to be learned by the learning agent. If the attacker wants to mislead the agent into certain bad behavior, it has to identify the set of policies that can be learned by the learner and can only choose bad behavior from that set.
  4. 4.
    Black box attack: As discussed previously, a practical attacker should be oblivious to the learning agent. Without knowing the details of the offline learning algorithm, it would be impossible to predict the exact influence of a reward perturbation on the learning result of the algorithm, which is required by a majority of poisoning attack strategies against RL.
Contributions: To address the challenges above, we adopt a basic assumption about efficient offline learning algorithms. Under the assumption, we adopt an attack framework where the attacker can approximate the learning outcome of an efficient learning algorithm while being oblivious to the learning algorithm or the training process. We derive theoretical insights on developing attacks that the attacker should make high-performing policies appear low-performing to the agent and vice versa. Based on the insights, we propose our main attack method: policy contrast attack, that can be used to assess the practical robustness of offline RL algorithms. To the best of our knowledge, this is the first black-box reward poisoning attack in the general offline RL setting. The core idea of our attack is to make the learner believe that the high-performing policies supported by the dataset are sub-optimal while some low-performing policies are optimal. We empirically test the efficiency of our attack on various standard datasets from the D4RL benchmark (Fu et al., 2020) learned by different state-of-the-art offline RL algorithms (Kumar et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2021). We find that our attack is efficient in a majority of cases with a limited attack budget. We also show that our attack is not sensitive to the choice of hyper-parameters. We hope that our work can inspire the development of more robust offline RL algorithms.
2 Preliminaries
Offline RL: We consider a standard offline RL setting (Cheng et al., 2022) where an agent trains on an offline dataset to learn how to perform well in an RL environment. An RL environment is characterized by an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}=\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition function, and $\mathcal{R}=\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is a mapping from the state space to a distribution over action space suggesting the way one behaves in the environment. Without loss of generality, we assume the initial state is always $s_{0}$. Then the performance of a policy $\pi$ for the environment is defined as $J_{\mathcal{R}}(\pi)=\mathbb{E}_{\pi,\mathcal{P}}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})|a_{t}\sim\pi(s_{t})]$, which represents the long-term discounted cumulative rewards for running the policy in the environment.
An offline reinforcement learning dataset is a collection of observation tuples collected from the environment: $\mathcal{D}=\{(s_{i},a_{i},s^{\prime}_{i},r_{i}),i=[0,\ldots,N]\}$ where $s^{\prime}_{i}\sim\mathcal{P}(s_{i},a_{i})$ and $r_{i}\sim\mathcal{R}(s_{i},a_{i})$. The state action distribution $\mu=\{(s_{i},a_{i}),i=[0,\ldots,N]\}$ represents the trajectories covered in the dataset. An offline reinforcement learning agent has access to the offline dataset $\mathcal{D}$ as well as the state and action spaces $\mathcal{S},\mathcal{A}$, and its goal is to find a high-performing policy.
Reward poisoning attack against offline RL: We consider a reward poisoning attack model where a malicious adversary can poison the reward signals in the offline dataset to mislead the learning agent. Formally, for the $i^{\text{th}}$ observation tuple $(s_{i},a_{i},s^{\prime}_{i},r_{i})$, the attacker can inject a perturbation $\Delta_{i}$ to the reward signal, then the corrupted observation tuple becomes $(s_{i},a_{i},s^{\prime}_{i},r_{i}+\Delta_{i})$. We denote $\mathbf{\Delta}={\Delta_{i},\ldots,\Delta_{N}}$ to be the corruption strategy of the attack.For a practical threat model, the attacker should have limited abilities. Formally, we consider the following constraints on the attacker.
  1. 1.
    No access to the offline training process: Since the training is offline, the learning agent can host the training process on local machines. Therefore, it would be impractical for the attacker to observe the training process of the learning agent.
  2. 2.
    No knowledge of the learning algorithm: There are a variety of efficient offline RL algorithms with very different learning strategies. Unless the learning agent announces its learning algorithm to the public, it would be impractical for the attacker to assume which learning algorithm the agent would choose. Therefore, the attacker should not have any detailed knowledge of the learning algorithm except for some basic assumptions that the algorithm is efficient when there is no attack.
  3. 3.
    Limited Budget: To make the attack stealthy, the attacker should make the corruption as small as possible. First, the attacker should limit the maximal perturbation on each reward signal $||\mathbf{\Delta}||_{\infty}=\max|\Delta^{i}|$, as an extreme value of reward would make the data point very suspicious. Second, the attacker should also limit the total amount of perturbation $||\mathbf{\Delta}||_{1}=\sum_{i=0}^{N}|\Delta^{i}|$. This is because poisoning the dataset would involve accessing the dataset collection process, and it will be suspicious if the attacker frequently accesses the process.
The general goal of the attacker is to make the learning agent learn a low-performing policy from the poisoned dataset. Formally, we adopt the following criteria for measuring the efficiency of an attack. Denote $B=||\mathbf{\Delta}||_{\infty}$ and $C=||\mathbf{\Delta}||_{1}$ to be the attack budgets required by the attack. Let $V=\mathbb{E}[J(\hat{\pi})]$ where $\hat{\pi}$ is the policy learned by the learning algorithm under the attack $\mathbf{\Delta}$, and the expectation is over the randomness in the learning process. An attack is more efficient if, with the same value of $B$ and $C$, it has a lower value of $V$ compared to other attacks.
3 Methods
3.1 Adversarial Reward Engineering
Here, we introduce the framework we adopt to solve the above attack problem. Given that the attacker is unaware of the learning agent, it is necessary to establish certain general assumptions that apply to the algorithms employed by the agent. Considering that a learning agent must ensure the efficiency of its learning algorithm when working with an uncorrupted dataset, it becomes practical for the attacker to make certain basic assumptions regarding the learning algorithm’s efficiency. One common characteristic of current offline RL algorithms is referred to as ’pessimism’ (Levine et al., 2020). In offline RL, a major challenge is known as ’distribution mismatch’. It says that if the distribution of trajectories generated by a policy is rarely covered by the dataset, the estimation of the policy’s performance based on that dataset will be inaccurate. Hence, the idea of ’pessimism’ aims to exclude policies that are inadequately covered by the dataset. To achieve this, state-of-the-art learning algorithms utilize diverse approaches and offer various theoretical guarantees under different assumptions (Kidambi et al., 2020; Fu et al., 2020; Fujimoto & Gu, 2021; Cheng et al., 2022; Wu et al., 2019). Consequently, exactly characterizing the learning efficiency of these algorithms with a universal mathematical formulation is difficult. Therefore, we adopt the following simple and fundamental assumption to assess the general efficiency of an offline RL algorithm
Assumption 3.1 ($\delta$-Efficient offline learning algorithm).
Given an offline RL dataset $\mathcal{D}$ collected from an environment $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$ with state-action distribution $\mu$. There exists a class of policies $\Pi_{\mu}$ that is supported by the distribution $\mu$. Let $\pi_{0}$ be the policy learned by the learning algorithm. For an efficient offline RL algorithm, there exists a small value $\delta>0$ such that $J_{\mathcal{R}}(\pi_{0})\geq\max_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)-\delta$. That is, the algorithm will learn the policy with nearly the highest performance on $\mathcal{M}$ among the supported policies.
For the rest of the section, we assume that the learning agents always use a $\delta$-efficient learning algorithm. Based on the assumption, similar to the previous work in the online learning setting (Xu et al., 2022), we can address the attack problem by solving a reward function engineering problem. We call the framework ‘adversarial reward engineering’, which is formally defined as below.
Definition 3.2.
Under the adversarial reward engineering framework, the attacker constructs an adversarial reward function $\widehat{\mathcal{R}}$. The corresponding corruption on the $i^{\text{th}}$ data $(s_{i},a_{i},r_{i})$ from $\mathcal{D}$ satisfies $\Delta_{i}=\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}$.
In the adversarial reward engineering framework, it is equivalent to say the corrupted dataset is collected from the adversarial environment ^$\widehat{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\widehat{\mathcal{R}})$ with the same state-action distribution $\mu$ as the original dataset. Under Assumption 3.1, Theorem 3.3 generally characterizes the efficiency of an attack under the adversarial reward engineering framework.
Theorem 3.3.
Under an attack within the adversarial reward engineering framework, any efficient learning algorithm satisfying Assumption 3.1 will learn the policy with a near-optimal performance on $\widehat{R}$: $J_{\widehat{R}}(\hat{\pi}^{*})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta$. The efficiency of the attack satisfies $V=J_{\mathcal{R}}(\hat{\pi}^{*})$, $B=\max_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$ and $C=\sum_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$
In general, the attacker needs to find a suitable $\widehat{\mathcal{R}}$ such that the required budget is limited, and the near-optimal policies in the adversarial reward function have poor performance in the true environment.
3.2 Inverted Reward Attack
Following the idea of reward engineering, one can immediately find an attack that can make the agent learn the worst possible policy under Assumption 3.1.
Definition 3.4 (Fully inverted reward attack).
The fully inverted reward attack sets $\widehat{\mathcal{R}}=-\mathcal{R}$ as the adversarial reward function where $\mathcal{R}$ is the true reward function of the underlying environment behind the dataset. The corresponding attack strategy of the fully inverted reward attack satisfies $\Delta_{i}=-2\cdot r_{i}$.
The inverted reward attack flips the sign of all the rewards in the dataset. In Theorem 3.5 we show the efficiency of the fully inverted reward attack.
Theorem 3.5.
Under the fully inverted reward attack, the learning agent will learn the policy supported by the dataset with nearly the worst performance $V\leq J_{\mathcal{R}}(\pi_{0})=\min_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)+\delta$. The budgets required by the attack are $B=2*\max_{i}|r_{i}|$ and $C=2*\sum_{i}|r_{i}|$.
Maximizing the cumulative reward on $-\mathcal{R}$ is equivalent to minimizing that on $\mathcal{R}$. Therefore, an efficient learning algorithm will learn the policy with the worst performance on the true reward function among the policies supported by the dataset. Despite achieving the worst outcome for the learner, the inverted reward attack faces the problem of requiring high budgets which is impractical. Since a practical attack should only corrupt a small portion of data, we consider an attack that can only randomly invert a part of the reward. We call this attack the ‘random inverted reward attack’. This attack strategy has been widely used in previous works as a non-trivial attack (Ye et al., 2023; Zhang et al., 2022).
Definition 3.6 (Random inverted reward attack).
The random inverted reward attack has a parameter $p\in(0,1)$. The attack randomly samples $p\cdot N$ indexes from $[1,\ldots,N]$ with replacement, where $N$ is the size of the dataset. Let the set of indexes be $I$. The attack strategy of the inverted reward attack satisfies $\Delta_{i\in I}=-2\times r_{i}$ and $\Delta_{i\not\in I}=0$.
We cannot find any strong guarantee for the effect of the attack on the learning agent. Empirically, we observe that when the attack inverts $<50\%$ of the states, the learner is usually still able to learn a policy almost as good as the one it learns in the uncorrupted dataset. Since the random inverted reward attack is not practical, we need to find a more efficient attack that can work with less budget yet still makes the agent learn a policy of low performance.
The inverted reward attack treats different state actions from the dataset equally, yet the reward for some state actions can have more influence on the learning process. Therefore, it is possible to construct more efficient attacks that focus the attack budget on the rewards in some specific state actions.
3.3 Policy Contrast Attack
Here we introduce our attack. First, in Theorem 3.7 we give a sufficient condition and another necessary for the adversarial reward function to make the agent learn a low-performing policy. The two conditions are the same when $\delta$ becomes 00.
Theorem 3.7.
Consider an attack in the adversarial reward engineering framework with $-^\widehat{\mathcal{R}}$. To make the actual performance of the learned policy lower than a value $J_{\mathcal{R}}(\pi_{0})<V$, a sufficient condition is that the adversarial reward function satisfies
$\exists\pi_{1}\in\Pi_{\mu}\text{. }J_{\mathcal{R}}(\pi_{1})<V,\forall\pi_{2}\in\Pi_{\mu}.J_{\mathcal{R}}(\pi_{2})\geq V,$
$J_{\widehat{\mathcal{R}}}(\pi_{1})>J_{\widehat{\mathcal{R}}}(\pi_{2})+\delta,$
and a necessary condition that it should satisfy is
$\exists\pi_{1}\in\Pi_{\mu}\text{. }J_{\mathcal{R}}(\pi_{1})<V,\forall\pi_{2}\in\Pi_{\mu}.J_{\mathcal{R}}(\pi_{2})\geq V,$
$J_{\widehat{\mathcal{R}}}(\pi_{1})>J_{\widehat{\mathcal{R}}}(\pi_{2})-\delta,$
In both cases, we call $\pi_{1}$ that satisfies the conditions ‘bad policy’ and $\pi_{2}$ ‘good policy’.
Theorem 3.7 suggests that the attacker should make a low-performing policy have a higher performance, in the adversarial reward function, than any high-performing policies. Formally, denote $\Delta J(\pi)=J_{\widehat{\mathcal{R}}}(\pi)-J_{\mathcal{R}}(\pi)$. Then to make $J_{\widehat{\mathcal{R}}}(\pi_{1})>J_{\widehat{\mathcal{R}}}(\pi_{2})+\delta$, the attacker needs to ensure that $\Delta J(\pi_{1})-\Delta J(\pi_{2})>J(\pi_{2})-J(\pi_{1})+\delta$. In other words, the attacker needs to make $\Delta J(\pi_{1})$ high for some bad policy $\pi_{1}$ and $\Delta J(\pi_{2})$ low for any good policy $\pi_{2}$. Inspired by this, at a high level, the key idea for our attack is that in the adversarial reward function, some low-performing policies appear as high-performing (bad policies look good), while any high-performing policy appears as low-performing (good policies look bad). Based on the idea, we construct our attack called ‘policy contrast attack’.
Bad policies look good: Following the idea, the first step of the attacker is to find some low-performing policies supported by the dataset. To achieve this, the attacker can train with any efficient algorithm on the dataset with the inverted reward. By Theorem 3.5, the attacker can learn the worst-performing policy supported by the dataset, which is exactly the goal. Intuitively, this policy and the policies that behave similarly to it should have low performance in practice. Therefore, to make some low-performing policies have a relatively high performance in an adversarial environment, a straightforward way is to increase the reward associated with the actions close to that given by the worst-performing policies. For this purpose, we denote the distance measure between two actions $a_{1},a_{2}$ as $d_{a}(a_{1},a_{2})$. For the continuous action space as the ones considered in the continuous control settings (Fu et al., 2020), we choose the distance measure as $d_{a}(a_{1},a_{2})=||a_{1}-a_{2}||_{2}$.
Good policies look bad: The next step of the attacker is to find the high-performing policies supported by the dataset and make them have low performance in the adversarial reward function. A straightforward idea is to find a set of different good policies such that any good policy will behave similarly to one of them. Following this idea, we use an iterative learning approach to find good policies. The high-level idea is to learn a high-performing policy different from previously learned policies at each iteration. The algorithm to find good policies is described in Alg 1.
Algorithm 1 Learning Set of Good Policies
  Input: dataset $\mathcal{D}$ of size $N$, number of iterations $K$, offline RL algorithm Alg
  Params: distance threshold $d$, corruption $\Delta$
  Initialize $\Pi_{2}=\emptyset$, $\widehat{\mathcal{D}}=\mathcal{D}$.
  for $i=1$ to $K$ do
     Learn $\pi$ from dataset $\widehat{\mathcal{D}}$
     Update $\Pi_{2}=\Pi_{2}\cup\{\pi\}$
     for $j=1$ to $N$ do
        Get state-action-reward $(s_{i},a_{i},r_{i})$ from $\widehat{\mathcal{D}}$
        if $\min_{\pi\in\Pi_{2}}d_{a}(a_{i},\pi(s))\leq d$ then
           Modify $r_{i}=-\Delta$
        end if
        Update $(s_{i},a_{i},r_{i})$ in $\widehat{\mathcal{D}}$
     end for
  end for
  Output: $\Pi_{2}$
In the iterative learning of good policies, the attacker starts with an empty set of good policies $\Pi_{2}=\emptyset$. At the first iteration, the attacker trains on the original dataset with an efficient learning algorithm and then adds the learned policy to the good policy set $\Pi_{2}$. In the subsequent iterations, the attacker modifies the dataset by decreasing the rewards at the actions close to one of the actions given by the policies from the good policy set $\Pi_{2}$. Note that this modification is not the corruption for the agent’s training. The attacker trains on the modified dataset and then adds the learned policy to the good policy set. Theorem 3.8 shows that policies of high performance will behave similarly to some policies from the good policy set.
Theorem 3.8.
Let $V$ be the performance of the learned good policy in the last iteration. For any policy $\pi\in\Pi_{\mu}$ such that $\min_{s\in\mathcal{S},\pi_{2}\in\Pi_{2}}d_{a}(\pi(s),\pi_{2}(s))>d$, its performance satisfies $J_{\mathcal{R}}(\pi)<V+\delta$.
To make the policies from the good policy set have bad performance in the adversarial reward, one can simply decrease the rewards associated with the actions close to the ones given by the good policies.
Algorithm 2 Policy Contrast Attack
  Input: dataset $\mathcal{D}$ of size $N$, efficient offline RL algorithm Alg
  Params: distance threshold $d$, corruption parameters $\Delta_{1}$, $\Delta_{2}$, number of iterations $K$
  Initialize $\widehat{\mathcal{D}}=\{\}$.
  Learn a bad policy $\pi_{1}$ from dataset with inverted rewards $-\mathcal{D}$
  Learn a good policy set $\Pi_{2}$ through Alg 1 with $K$ iterations
  for $i=1$ to $N$ do
     Get state-action-reward $(s_{i},a_{i},r_{i})$ from $\mathcal{D}$
     if $\min_{\pi\in\Pi_{2}}d_{a}(a_{i},\pi(s_{i}))\leq d$ then
        Modify $\hat{r}_{i}=-\Delta_{2}$
     end if
     if $d_{a}(a_{i},\pi_{1}(s_{i})\leq d$ then
        Modify $\hat{r_{i}}=r_{i}+\Delta_{1}\cdot(1-d_{a}(a_{i},\pi_{1}(s_{i})/d)$
     end if
     Update $\widehat{\mathcal{D}}=\widehat{\mathcal{D}}\cup\{(s_{i},a_{i},\hat{r}_{i})\}$.
  end for
  Output: $\widehat{\mathcal{D}}$
Figure 1: Peformance of learning algorithms on different datasets under the attacks. We take the title of the first figure as an example to explain the meaning of the title. ‘HalfCheetah’ means the RL environment is HalfCheetah; ’Medium Expert’ means the dataset is collected by a mixture of medium and expert policies. ‘IQL attacks TD3_BC’ means the learning algorithm used by the attacker is IQL, and the one used by the learning agent is TD3_BC.
Practical implementation: Finally, we consider the practical implementation of the attack method. First, the attacker finds a bad policy $\pi_{1}$ by training on the dataset with inverted rewards and a good policy set $\Pi_{2}$ by Alg 1. Intuitively, with a higher value of $K$ for the size of $\Pi_{2}$, more good policies of different behaviors will be covered. Empirically, we show that the performance of the attack is not sensitive to the value of $K$. Then, the attacker modifies the reward through an adversarial reward function that increases the rewards for bad actions and decreases the rewards for good actions. In Alg 2 we show the pseudo code for the policy contrast attack. The adversarial reward function in Alg 2 satisfies:
^$\widehat{\mathcal{R}}(s,a)=\begin{cases}\mathcal{R}(s,a)+\Delta_{1}\cdot\\ (1-\frac{d_{a}(a,\pi_{1}(s))}{d}),&\text{if }d_{a}(a,\pi_{1}(s_{2}))\leq d\\ -\Delta_{2},&\text{else if }\\ &\min_{\pi_{2}\in\Pi_{2}}d_{a}(a,\pi_{2}(s))\leq d\\ \mathcal{R}(s,a),&\text{otherwise}\end{cases}$
Here $\Delta_{1}$, $\Delta_{2}$, and $d$ are the hyper-parameters of the attack determined by the attack budgets. The first part of the reward function is to increase the rewards associated with the bad actions given by the bad policy. The formation of the reward modification in this part follows Xu & Singh (2023). The second part of the reward function is to decrease the rewards associated with good policies.
Note that there are some cases where the actions given by the good policies and the bad policy are close. In this case, there will be a conflict for the aforementioned corruption strategy. Intuitively, these cases are rare as usually good and bad actions should not be similar. In addition, in these cases, a change in reward has a similar influence on the performances of good and bad policies in the adversarial function, suggesting that the corruption in these cases has a limited impact on the ranking of good and bad policies in the adversarial reward. To ensure that the bad policy always has a higher performance in the adversarial environment, we always increase the reward associated with the bad actions whenever there is a conflict.
At last, we want to discuss why the policy contrast attack should be more efficient than the random inverted attack. At a high level, both the policy contrast and the inverted reward attack want to inverse the agent’s ranking of good and bad policies. In contrast to the inverted reward attack that treats all state actions equally and inverse the performance of all policies, the policy contrast attack focuses on policies that matter more as pointed out by Theorem 3.7 and the state actions that are more important. This potentially makes the policy contrast attack more efficient in terms of budget.
Figure 2: Influence of different $B$ budget on the attack.
Figure 3: Influence of different $C$ budget on the attack.
4 Experiments
4.1 Performance of Different Attack Strategy
We test the effectiveness of our attack against current state-of-the-art offline RL algorithms training on the standard D4RL offline RL benchmark’s continuous control domains (Fu et al., 2020). we show that with limited attack budgets, our policy contrast attack is in general more effective.
Experiment setup: The offline datasets we choose from the D4RL benchmark include different combinations of environments and types of trajectories. We consider HalfCheetah, Hopper, Walker2d, and Ant environments. We consider the data collected by a single medium policy (medium), a single expert policy (expert), a single random policy (random), a mixture of medium and expert policies (medium-expert), and the replay buffer of a medium policy (medium-replay).
For the learning algorithms, we choose IQL  (Kostrikov et al., 2021), CQL (Kumar et al., 2020), and TD3-BC (Fujimoto & Gu, 2021), which are frequently used as baselines in current offline RL studies. The implementation of these algorithms is based on CORL library (Tarasov et al., 2022). Recall that the attacker also needs a learning algorithm to learn good and bad policies. To ensure that the attacker has no information about the learning agent, the learning agent and the attacker always use different learning algorithms.
For the baseline attack, we choose the random invert attack as it is the only non-trivial in the current literature. In the ablation study, we also show the result of using only one part of the attack strategy from the policy contrast attack, that is, the attack only makes good policy look bad or bad policy look good.
For the attack budget, following Ye et al. (2023), we always set the budget for $B$ and $C$ to be the amount required by the random inverted reward attack that corrupts $30\%$ of the data. We test the influence of different attack budgets on the efficiency of the attack in the ablation study.
For the hyper-parameters of the attack, the corruption parameters $\Delta_{1}$ and $\Delta_{2}$ are set to maximally utilize the budget $B$, and the attack radius $d$ is set to maximally utilize the budget $C$. The number of good policies, which is also the number of iterations $K$, is set as $5$ here. In the ablation study, we show that the efficiency of the algorithm is not sensitive to the choice of corruption parameters and the number of good policies. For each experiment, we run $5$ times with different random seeds and show the average result.
Results: The offline RL algorithms we choose iteratively update their learned policies. To intuitively show the details of their training process under the attacks, we evaluate and show the performance of the policy learned by the algorithm at each iteration during training. In the appendix, we show the quantified result of the learning algorithms’ performance under the attack.
Fig 1 shows the main results of running our policy contrast attack in different learning scenarios. The naive baseline is the training with no attack. The ideal effect of the attack is the training with fully inverted reward as the learning algorithm tries to learn the worst-performing policy in this case. We observe that our policy contrast attack is usually effective for any combination of datasets and learning algorithms while the random inverted attack has almost no effect on the learning result except for the Hopper-Medium Replay dataset where both attacks are similarly effective.
Figure 4: Influence of different sizes of good polices on the attack.
Figure 5: Attacks that only make bad policies look good and good policies look bad.
4.2 Ablation Study
Attack budget $B$: Here we study the influence of per-step attack budget $B$ on the attack efficiency. Denote $r_{\text{max}}=\max_{i}|r_{i}|$ be the maximal absolute reward from the dataset. In the main results, the budget required by the attacks is $B=2*r_{\text{max}}$. Here we set the value of $B$ to be $1.5\cdot r_{\text{max}}$ and $1\cdot r_{\text{max}}$ and test on two randomly chosen datasets. The results in Fig 2 show that our policy contrast attack remains efficient with a lower budget of $B$. We note that the attack is slightly more efficient with less value of $B$. Intuitively, although the per-step corruption is less, the attack influences the performance of more policies, making it possible for the agent to learn slightly worse policies. The result also suggests that the efficiency of our attack is not significantly affected by the choice of the corruption parameters $\Delta_{1}$ and $\Delta_{2}$ at each step as their values depend on $B$.
Attack budget $C$: Here we study the influence of attack budget $C$, the total amount of corruption, on the efficiency of the attack. We set the value of $C$ to be the amount required by a random inverted attack that corrupts $20\%$ and $40\%$ of data. The results in Fig 3 show that our attack is more efficient with a higher budget of $C$, and it remains effective with a lower budget of $C$.
Size of good policy set: Here we study the influence of the size of $\Pi_{2}$ on the efficiency of the attack. We set the sizes of the good policy set to be $5,7,10$. In Fig 4, we observe that the learning outcomes of the algorithms are similar for different sizes of good policy sets. The results suggest that our policy contrast attack is not considerably affected by the number of good policies.
Only Good/Bad Policies attacks: To show that it is beneficial to do both making good policies look bad and bad policies look good, we test the efficiency of the attack strategy with only one part. That is, the attack either only decreases the reward associated with the good policies or increases the reward associated with the bad policy. In Fig 5, we observe that both attacks lose some efficiency compared to the policy contrast attack. Here we provide empirical insights on why these two attacks do not work alone. For making good policies look bad, it can happen in practice that the good policy set does not cover all policies of high performance, therefore the agent is still able to identify a high-performing policy. For making a bad policy look good, in practice even if the bad policy already has the highest performance in the adversarial environment, without the good policy attack part, the learning agent may not be able to learn it and still converge to a good policy as the adversarial reward function is more complicated. In this case, by breaking the optimality of the good policies, it becomes easier for the learning agent to converge to the bad policy.
5 Related Work
Reward poisoning attack in online RL: Most studies on reward poisoning attacks against RL focus on the online learning setting. Rakhsha et al. (2020); Zhang et al. (2020) theoretically investigate the attack problem in the online RL setting with tabular MDPs. They show that given a specific learning algorithm or an algorithm with certain learning guarantees, attack strategies exist such that the learning algorithm will learn a low-performing policy even if the attack only slightly perturbed the training process.
In the more complicated general RL setting where the learning agent needs deep neural networks for function approximation, Sun et al. (2020) empirically show that current state-of-the-art DRL algorithms are vulnerable under the poisoning attack when the attacker is aware of the learning algorithm. Xu et al. (2022); Xu & Singh (2023) further shows that even in the black-box setting where the attacker is oblivious to both the environment and the learning algorithm, the data poisoning attack can achieve both targeted and untargeted attack goals with limited budgets.
Reward poisoning attack and defense in offline RL: There is very limited work on the poisoning attack in offline RL settings. Ma et al. (2019) theoretically studies the problem for tabular MDPs. For the defense against the attack, Zhang et al. (2022); Ye et al. (2023) design algorithms and theoretically prove their robustness. We show that our attack is much more efficient than the random attack used for empirical evaluation in Ye et al. (2023).
6 Conclusion
In this work, we propose the first black-box reward poisoning attack in offline reinforcement learning. We design an attack strategy based on our novel theoretical insights. Our empirical results suggest that the attack is effective in various learning scenarios. The limitation of our work is that we only consider black-box attacks, and the attacker requires access to the full offline training dataset.
7 Impact and Ethics
This paper presents work whose goal is to advance the field of Machine Learning. Our proposed methods expose vulnerabilities of existing offline DRL methods which can be exploited by adversaries. We hope that the issues highlighted in our work will motivate the development of more robust offline DRL algorithms.
Appendix A Proof for Theorems and Lemmas
Proof for Theorem 3.3 By the definition of the attack, the perturbation on the $i^{\text{th}}$ data is $\Delta_{i}=\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}$. Therefore, the total corruption required by the attack is $C=\sum_{i}|\Delta_{i}|=\sum_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$, and the per-step corruption is $B=\max_{i}|\Delta_{i}|=\max_{i}|\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}|$. By Assumption 3.1, the agent should learn a near-optimal policy on the reward function of the dataset, which is $\widehat{\mathcal{R}}$ under the attack. Therefore, the policy $\hat{\pi}^{*}$ learned by the agent under the attack satisfies $J_{\widehat{R}}(\hat{\pi}^{*})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta$.
Proof for Theorem 3.5 The adversarial reward function constructed by the fully inverted attack is $\widehat{\mathcal{R}}=-\mathcal{R}$. By Assumption 3.1, the agent will learn a policy $\pi_{0}$ that is near-optimal on $\widehat{\mathcal{R}}$: $J_{\widehat{\mathcal{R}}}(\pi_{0})\geq\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta$. Therefore, the performance of the learned policy on the true reward function satisfies $J_{\mathcal{R}}(\pi_{0})=-J_{\widehat{\mathcal{R}}}(\pi)\leq-(\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)-\delta)=\min_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)+\delta$.
Proof for Theorem 3.7 First, we prove the sufficient condition. For any policy $\pi_{2}\in\Pi_{\mu}:J_{\mathcal{R}}(\pi_{2})>V$, we have $\max_{\pi\in\Pi_{\mu}}J_{\widehat{\mathcal{R}}}(\pi)\geq J_{\widehat{\mathcal{R}}}(\pi)\geq J_{\widehat{\mathcal{R}}}(\pi_{2})+\delta$. Therefore, $\pi_{2}$ is not a near-optimal policy on the adversarial reward function, and by Assumption 3.1, the agent will not learn such a policy.
Next, we prove the necessary condition. Let $\pi_{0}$ be the policy learned by the agent under the attack, and this policy has a performance less than $V$ on the actual reward function $J_{\mathcal{R}}(\pi_{0})<V$. Therefore, we can take $\pi_{1}=\pi_{0}$. By Assumption 3.1, the gap between the performance of $\pi_{1}$ and any other policy on $\widehat{\mathcal{R}}$ must be less than $\delta$. Therefore, for any policy $\pi_{2}\in\Pi_{\mu}:J_{\mathcal{R}}(\pi_{2})>V$, we have $J_{\widehat{\mathcal{R}}}(\pi_{1})>J_{\widehat{\mathcal{R}}}(\pi_{2})-\delta$.
Proof for Theorem 3.8 At the last iteration of Alg 1, the performance of the policy $\pi_{K}$ learned in the iteration on the adversarial reward function ^$\widehat{\mathcal{R}}_{K}$ in the iteration is no greater than $V$ because the adversarial reward function is strictly less than the actual reward function. By Assumption 3.1, the performance of any policy $\pi\in\Pi_{\mu}$ on ^$\widehat{\mathcal{R}}_{K}$ satisfies $J_{\widehat{\mathcal{R}}_{K}}(\pi)\leq V+\delta$. For any policy $\pi\in\Pi_{\mu}:\min_{\pi_{2}\in\Pi_{2},s}||\pi(s)-\pi_{2}(s)||_{2}>d$, their performance on the actual reward function and the adversarial reward function are the same. Therefore, their actual performance must be less than $V+\delta$.
Appendix B Quantified Experiment Results
Here we show the quantified learning performance of the learning algorithms training on the dataset under the attacks. The score we show in Table 1 is the D4RL normalized score for the performance of the optimal policy learned in the last $10$ checkpoints during training.
D4RL Dataset
Normal Training
Random 30% Invert
Policy Contrast Attack
Full Invert
HalfCheetah-Medium-Expert-v2
91.03
89.83
76.23
13.10
Hopper-Medium-Expert-v2
105.89
99.87
82.02
16.87
HalfCheetah-Expert-v2
95.50
94.9
93.84
80.88
Walker2D-Medium-v2
81.84
77.38
73.10
67.66
Walker2D-Medium-Replay
80.17
78.74
31.85
-0.22
Hopper-Medium-Replay-v2
85.20
25.14
30.27
0.65
Table 1: Quantified performance of learning algorithms under the attacks.